# The influence of the 3D Galactic gas structure on cosmic-ray transport and -ray emission

### v1

@@ 3 Results and Discussion
The main results of our simulations are the energy-dependent spatial distributions of Galactic CR fluxes () and of local -ray emissivities () ā as well as the LOS integrals of the latter ā computed with the 3D model. In this section, we present the summary statistics of these results and compare them, where applicable, to the morphology of the gas reconstruction maps and the results obtained by using the 2D model. All uncertainties shown here correspond to the 1- region of the simulation outputs. For all our results related to CR nuclei fluxes we define the energy axis by their corresponding kinetic energy per nucleon (n). We denote this convention as in GeV/n units. For leptons, we use their corresponding kinectic energy in GeV units.
@@ 3.1 Impact on Transport Parameters
Boron nuclei are predominantly produced via the interaction of heavier nuclei (e.g. Carbon and oxygen) with ISM gas. The relative abundances between secondary and primary nuclei are sensitive to the mechanisms of CR transport, making their measurement an important probe to constrain the parameters involved in CR propagation (Strong etĀ al., 2007b; DAMPE Collaboration, 2022). Particularly, Boron and Carbon abundances are extensively well documented. Therefore, we use the available B/C ratio data to investigate the constraints introduced in the transport parameters related to diffusion, given the usage of the gas reconstructions. In Fig 2 we show the fits to the B/C ratio data resulting from the calibration of the transport parameters described in section 2. We note that the 2D model requires and values slightly above typical values (see e.g., Strong etĀ al., 2007b, where ām2ās-1 and ākmās-1).
@@
The transport parameters of the 3D model deviate significantly from the ones in the 2D model by requiring a smaller diffusive strength to remain close to the AMS-02 data. This behaviour is related to the gas density near Earth in both models. We check this behavior in each model by evaluating the corresponding average hydrogen densities within a sphere of 4kpc radius around the location of Earth. For the 2D model, ācm-3 and for the 3D model, ācm-3. We find that in this region, leading to and . Consequently, we observe a direct correlation between the average gas density in the vicinity of Earth and with and , where smaller average gas densities require weaker diffusion and reacceleration. However, we note that changes in do not translate linearly to changes in and , this results from gas structures at different scales affecting the flux of CRs via their interactions (e.g. spallation, energy losses), which we discuss further in section 3.3.
@@
In Fig 3, we present the simulated Carbon and Boron spectra at Earth that constitute the fitted B/C ratio, as well as corresponding observations. The main deviations between data and simulated spectra are present for energies below 10āGeV/n, where at the estimated heliospheric modulation, the simulated fluxes are below the data. Stable Boron isotopes (i.e. 10B and 11B) are mainly produced from spallations of CR 12C and of CR 16O with ISM protons, which also produce unstable Carbon isotopes that decay further (see Table 1 in Heinbach & Simon (1995) for the related cross-sections). This anti-correlates the population of Boron and Carbon non-linearly, where the latter is dominated by the 12C (see e.g. Fig 6). As a result, given enough local gas density, we expect a greater gain of Boron nucleiisotopes from a lower loss of Carbon nuclei., Twhich can explain therefore, the relative deviations from the data beingare larger for Carbon than for Boron result from the sensitivity of the distribution of CRs to both the choice of transport parameters and the local gas density. We also note that for these models ā in order to account to the change in slope to fit AMS-02 data ā the diffusion coefficient indices and deviate from the standard Kolmogorov spectrum index .
@@ 3.2 Impact on CR distributions
All nuclei are subject to spallation in dense gas regions during transport, resulting in the production of secondary lighter nuclei at the expense of the interacting primary. As a consequence of this effect, simulations display under-populated regions that appear for primary nuclei where high density gas regions are located. In contrast, for secondaries over-populated regions appear at high gas density regions. Hence, particles of mainly primary (secondary) nature are anti-correlated (correlated) with the distribution of gas. We illustrate this morphological impact of the 3D model on the distribution of different CR species. For this purpose, we look at the flux distributions at 10āGeV/n and 1āTeV/n of: Carbon 12 (12C), Boron 10 (10B), primary and secondary protons, and primary and secondary electrons (see sections 3.2.1, 3.2.2 and 3.2.3, respectively).
To compare the differences between 2D and 3D models across all energies, we further investigate the CR spectra from the 2D model and the mean CR spectra from the 3D model, for the same species as listed before. For all spectra shown henceforth, we use the set of transport parameters given in Table 1 for the 3D model.
We calculate spectra at the Galactic center (GC) and at Earthās location, where the latter is fixed by the underlying velocity field model used to reconstruct the 3D gas density. For the BEG03 model, Earth is assumed to be located at 8.0ākpc from the GC. We choose the GC region to study the impact of the distribution of gas structures, given that the average gas density within any distance from the GC is the same between the 2D and 3D models. For spectra at Earth we compare with CR data, where we use the force-field approximation to model heliospheric modulation. We take the modulation potential (see e.g., Ghelfi etĀ al., 2017, for details on its estimation) from the AMS-02 measurements, as reported in the cosmic ray database (CRDB) (Maurin etĀ al., 2014, 2020, 2023).
3.2.1 Boron and Carbon
In Fig 4, we show the mean and uncertainty of the distribution of 12C and 10B fluxes in the Galactic plane, at 9.64āGeV/n and 1.04āTeV/n energies. For 12C, being mainly a primary CR particle, we note that under-populated regions are correlated with dense gas regions. This effect is most notable at 9.64āGeV/n near the GC, where the under-populated region follows the same profile as the dense gas region seen at the same location in Fig 1 for the 3D model. Furthermore, the under-populated regions disappear with increasing energy, becoming barely visible at 1.04āTeV/n and staying only at the densest gas regions. This result depends on the CR species and is a consequence of the corresponding dominant interaction at the energy of interest. When energy increases, loss time scales (e.g. for spallation, energy losses) become much longer when compared to the escape time of CRs from the Galaxy due to diffusion, causing spatial structures to become smooth.
For 10B, being of purely secondary nature, we note that the over-populated regions are directly correlated to dense gas regions. In contrast to 12C, this correlation is not suppressed with increasing energies, despite diffusion dominating over other interactions. This is a consequence of the secondary CR nuclei production rate being proportional to the ISM gas density. Therefore, given that losses are also high in the high density regions, secondaries can not easily escape these regions by diffusion. We discuss the relationship between gas and interaction rates further in section 3.3.
@@ 3.2.1 Boron and Carbon
For the corresponding uncertainties, we observe that they follow a profile similar to the reconstruction uncertainty of the gas within 10 kpc from the GC, as shown in Fig 1. The gas reconstruction profile is more evident at 1.04āTeV/n, whereas at 9.64āGeV/n and more clearly for 10B, structures towards the GC are smeared out and do not resemble the gas reconstruction profile. Despite this, for both species, we note that the relative uncertainties are much smaller for the CR distributions than for the gas reconstruction, decreasing further as energy increases.
In Fig 5 we present the B/C ratio at Earth and at the GC for both the 2D and 3D models using the same set of transport parameters (i.e. those from the 3D model, as discussed above). At both locations, we observe a higher B/C ratio in the 2D model than in the 3D model, where we find a general over-population of Boron at all energies. We investigate the effects that lead to this increase from the individual Carbon and Boron spectra at Earth and at the GC, which we show in Fig 6. At Earth, the total Carbon spectrum for the 2D model shows a slight decrease below 30āGeV/n, converging to the same flux as in the 3D model above this energy. A slight hardening of the Boron spectrum is present in the 2D model above 10āGeV/n, while both models agree below this energy. This leads to the overall increase of the B/C ratio in the 2D model seen in Fig 5, when compared to the 3D model. At the GC, the Carbon flux for the 2D model is noticeably reduced up to an energy of 200āGeV/n with respect to the 3D model. For Boron, the 2D model exhibits a reduced flux up to an energy of 10āGeV/n with respect to the 3D case. At 10āGeV/n, the Boron spectra are harder in the 2D model, leading to more high-energy Boron produced at the GC than in the 3D model.
In general, we observe a common set of effects. All particles ā albeit more strongly for primaries than for secondaries ā possess a higher flux in the 3D model up to a certain energy. Secondary-particle fluxes are lower in the 3D model above a certain energy. The energy at which these fluxes change depends on the corresponding CR species, and the magnitude of these differences between models depends on the location, being more obvious when looking at locally dense gas regions. Therefore, the flux differences reflect the correlation between the distribution of gas with the mean-free-path a particle traverses, given the presence of local gas over-/under-densities through the particular path and sufficiently short interaction rates (see e.g., section 3.3 and Fig.Ā 13 for the computation of this correlation). In particular for Boron and Carbon in the 3D model, this means that a higher flux of low-energy primaries is correlated to a lower flux of high-energy secondaries, since less spallation processes take place due to the local gas under-densities of a structured 3D model.
The dependence on the distribution of gas can be studied further by examining the spectra at the GC in Fig, 6. Here, the gas density is locally the same in both the 2D and 3D model. Therefore, loss processes contribute in the same way and the resulting differences in fluxes between models come as a direct consequence of the distribution of gas structures itself, instead of the local average density. The differences are especially visible for the total Carbon spectra at energies below 10āGeV/n, where the flux for the 2D model is about 2 times smaller than the one for the 3D model. From Fig 1, we note that within 4ākpc of the GC in the map for the 2D model there is an under-dense ring to which CRs can diffuse with equal probability in any direction, whereas in the 3D model the local under-dense regions are located north and south of the GC, constraining the directions in which CRs can diffuse from the GC.
3.2.2 Primary and secondary protons
In Fig 7 we show the mean and uncertainty of the distribution of simulated protons in the Galactic plane where, similarly to the 12C and 10B cases, we find that the same correlation between primaries, secondaries and gas exists. In contrast, the distributions exhibit a different energy dependence, with far less structures visible for primary protons. As for the distribution of secondaries, while still correlated to the gas density, it is less saturated at lower energies for protons. This is a consequence of the proton interactions (see e.g., Tan & Ng, 1983) for the modeling of the relevant proton cross-section) being inherently different from the Carbon and Boron ones, making the imprint of the gas distribution on CR fluxes dependending also on the corresponding CR species.
In Fig 8 we show the simulated proton spectra, where we note again the general effects discussed for Carbon and Boron. Here, the differences between models are negligible at Earth, which comes as a consequence of primary protons dominating the total flux, and being used for the normalization of the fluxes of all nuclei. With the differences between models being more noticeable at the GC than at Earth, and at energies where catastrophic and momentum losses in gas regions are significant, we observe again the structure-dependent behaviour of the simulated CR fluxes. We note as well that this dependence is only broken when diffusion dominates at high energies.
3.2.3 Primary and secondary electrons
The relationship between electrons and gas density differs from the CR species considered so far, due to their different energy loss processes. One important process is bremsstrahlung, which takes place in dense gas and can dominate electron losses between hundred MeV and a few GeV, before becoming subdominant to IC losses due to interactions with the different components of interstellar radiation field (ISRF) (Evoli etĀ al., 2020). As for the production of secondaries, this occurs via the decay of mesons produced in collisions of CR nuclei with gas, producing secondary electrons in dense gas regions. In Fig.Ā 9 we show the mean distribution and standard deviation of the simulated electrons in the Galactic plane, where we see the same correlation structure discussed so far due to the interactions discussed before. Primary electron under-densities remain visible for the densest gas clouds at energies of up to 150āGeV. However, as energy increases losses due to IC and synchroton emission become dominant, and no further gas structures are imprinted on primaries. Consequently, given the high energy losses at TeV energies, very high energy electrons seem to not travel far from their sources ā or gas for secondary electrons ā (see e.g., Mertsch, 2020; Evoli etĀ al., 2021; Thaler etĀ al., 2023), leading to a sharper outline of their structures with increasing energies, as seen at 1.04āTeV in comparison to 9.64āGeV.
We show the electron spectra at Earth and at the GC in Fig 10. Once again, we note the general effects pointed out for Carbon and Boron. Like for the primary proton case, lepton species are normalized using primary electrons, causing negligible differences between models at Earth. Furthermore, the discrepancy to AMS-02 data can be addressed by adapting the injection spectra of electrons that were used in the simulations. Particularly in the electron case, secondary species dominate the spectrum at very high energies, due to the softening of the primary injection spectrum. We see this effect clearly at āTeV at the GC.
@@ 3.3 Correlation Structure
Here, we display the correlation structure observed thus far in detail, by focusing on CR protons first to discuss how gas structures modulate their energy-dependent morphology. For this purpose, we study the proton interaction timescales at both the GC and Earth, the low-energy mean distribution of CR protons, and the correlation between gas and CR protons in the 3D model. As previously stated, energy-dependent spatial structures follow from the spatial distribution of gas via the corresponding interactions of the different particle species. When such interactions contribute significantly to CR losses in comparison to the escape rate of CRs from the Galaxy, the resulting CR distribution exhibits local structures that follow the gas maps. In Fig. 11, we show a comparison for protons between their inelastic, ionization and escape loss rates averaged over a sphere with radius 4ākpc around Earth. The escape rate is computed as with ākpc the halo height and the rigidity-dependent diffusion coefficient, using the values for the 3D model in Table 1. To illustrate the impact of the gas distribution on the distribution of CR fluxes at lower energies, we show in Fig. 12 the Galactic plane distribution of primary and secondary protons at 108āMeV/n. Although ionization rates dominate the low energy regime, primary protons are still subject to inelastic losses, which in turn drive the production of secondaries. We note that the ionization losses dominate for energies lower than āMeV/n, resulting in decreased primary proton fluxes in most dense gas regions within 10ākpc of the GC. The under-dense features disappear at locations with sufficiently less dense gas as diffusion becomes dominant, leading to the smoothing of the distribution of primary protons visible in Fig. 7 for the distributions 9.64āGeV/n and 1.04āTeV/n. We note that the gas-induced structures that are still visible at 9.64āGeV/n result due to inelastic losses being on a similar magnitude as diffusion. For secondaries, while the loss rates are the same, their injection occurs only at dense gas locations and thus their distribution is primarily determined by the gas maps.
@@
To quantify this energy-dependent relationship between the structure of fluxes for primaries and secondaries, and that of the gas density, we calculate the Pearson correlation coefficient. For this, we use protons and electrons, and compare their distribution with the gas distribution inside 2 regions of radius ākpc and ākpc around the GC in the Galactic plane. The ākpc region considers the small over-dense patch of gas at the very center of the Galaxy, while the ākpc region takes into account local over- and under-dense gas clouds. The resulting energy-dependent correlation coefficients, along with the corresponding regions are shown in Fig. 13. We note the anti-correlation (correlation) between primaries (secondaries) and gas density in both regions at all energies. For very-high-energy secondary electrons (i.e. āTeV/n) the correlation is very close to 1, reflecting an almost linear dependence between gas and secondary electrons as a consequence of their increasingly strong confinement to their source regions (i.e. the high density regions) due to their high energy losses. We observe that with increasing energies, the absolute value of the correlation coefficient between secondaries and gas is bigger than the one between primaries and gas. This comes as a consequence of the anti-correlation with primaries being driven by gas-dependent losses. On the ākpc region, the correlation coefficient for secondary protons becomes considerably smaller, since they can travel larger distances than electrons due to lower energy losses.

### v2

@@ 3 Results and Discussion
The main results of our simulations are the energy-dependent spatial distributions of Galactic CR fluxes () and of local -ray emissivities () – as well as the LOS integrals of the latter – computed with the 3D model. In this section, we present the summary statistics of these results and compare them, where applicable, to the morphology of the gas reconstruction maps and the results obtained by using the 2D model. All uncertainties shown here correspond to the 1- region of the simulation outputs. For all our results related to CR nuclei fluxes we define the energy axis by their corresponding kinetic energy per nucleon (n). We denote this convention as in GeV/n units. For leptons, we use their corresponding kinectic energy in GeV units.
@@ 3.1 Impact on Transport Parameters
Boron nuclei are predominantly produced via the interaction of heavier nuclei (e.g. Carbon and oxygen) with ISM gas. The relative abundances between secondary and primary nuclei are sensitive to the mechanisms of CR transport, making their measurement an important probe to constrain the parameters involved in CR propagation (Strong et al., 2007; DAMPE Collaboration, 2022). Particularly, carbon and boron abundances are extensively well documented. Therefore, we use the available B/C ratio data to investigate the constraints introduced in the transport parameters related to diffusion, given the usage of the gas reconstructions. In Fig 2 we show the fits to the B/C ratio data resulting from the calibration of the transport parameters described in section 2. We quantify our goodness of fit for energies above 10 GeV to avoid the impact of heliospheric modulation, which we only consider via the force-field approximation. We find a value of 1.47. We note that the 2D model requires and values slightly below those used for the Y4201505 model in Ackermann et al. (2012) where m2 s-1 and km s-1.
@@
The transport parameters of the 3D model deviate significantly from the ones in the 2D model by requiring a smaller diffusive strength to remain close to the AMS-02 data. This behaviour is related to the gas density near Earth in both models. We check this behavior in each model by evaluating the corresponding average gas densities within a sphere of 4kpc radius around the location of Earth. For the 2D model, cm-3 and for the 3D model, cm-3. We find that in this region, leading to and . Consequently, we observe a direct correlation between the average gas density in the vicinity of Earth and with and , where smaller average gas densities require weaker diffusion and reacceleration. However, we note that changes in do not translate linearly to changes in and , this results from gas structures at different scales affecting the flux of CRs via their interactions (e.g. spallation, energy losses), which we discuss further in Section 3.3. We also note that in both models the diffusion coefficient indices and differ (i.e. ) and deviate from the standard dependence for Kolmogorov turbulence where . Similarly to and , the energy dependence of the cross sections used to parameterize the CR spallation and energy losses, when folded with either gas model, leads to a change in the slope of the simulated CR spectra. We address such a slope change by adjusting and to fit our simulated B / C ratio to the AMS-02 data. As discussed above, in the 4kpc spherical region centered on Earth, since , we note that the quantity is anti-correlated to . Furthermore, the distribution of local gas clouds – even for the same local among gas realizations – also affects the simulated CR flux, indirectly impacting the choice of transport parameters. We discuss this relation in more detail in section 3.2.
@@
In Fig 3, we present the simulated carbon and boron spectra at Earth that constitute the fitted B/C ratio, as well as corresponding observations. The main deviations between data and simulated spectra are present for energies below 10 GeV/n, where at the estimated heliospheric modulation, the simulated fluxes are below the data. Stable Boron isotopes (i.e. 10B and 11B) are mainly produced from spallations of CR 12C and of CR 16O with ISM protons, which also produce unstable Carbon isotopes that decay further (see Table 1 in Heinbach and Simon (1995) for the related cross-sections). This anti-correlates the population of carbon and boron non-linearly, where the latter is dominated by the 12C (see e.g. Figs 16 and 17). As a result, given enough local gas density, we expect a greater gain of Boron nuclei from a lower loss of Carbon nuclei. Therefore, the relative deviations from the data being larger for Carbon than for Boron result from the sensitivity of the distribution of CRs to both the choice of transport parameters and the local gas density.
@@ 3.2 Impact on CR distributions
All nuclei are subject to spallation in dense gas regions during transport, resulting in the production of secondary lighter nuclei at the expense of the interacting primary. As a consequence of this effect, simulations display under-populated regions that appear for primary nuclei where high density gas regions are located. In contrast, for secondaries, over-populated regions appear at high gas density regions. Hence, particles of mainly primary (secondary) nature are anti-correlated (correlated) with the distribution of gas. We illustrate this morphological impact of the 3D model on the distribution of different CR species, particularly of leptons and heavy and light nuclei, due to the different dominant interaction processes with ISM matter at different energies. For this purpose, we look at the distributions at 10 GeV/n and 1 TeV/n of: Carbon 12 (12C), Boron 10 (10B), primary and secondary protons, and primary and secondary electrons. (See Figs 4, 5 and 6, respectively for the corresponding mean CR flux distributions and uncertainties in the Galactic plane).
@@
We find that the corresponding CR uncertainties follow a profile similar to the gas uncertainty within 10 kpc from the GC, shown in Fig 1. For all species, the imprint of the morphology of the gas is evident for all energies. This imprint is stronger at lower energies, as seen at 9.64 GeV/n in contrast to 1.04 TeV/n. Particularly for secondaries (10B and secondary protons) at 9.64 GeV/n, structures do not fully resemble the gas reconstruction profile. We note that for both primary and secondary species the relative uncertainties are much smaller for the CR distributions than for the gas reconstruction, decreasing further with increasing energy.
We observe that under-populated regions of particles are present for all primary species and are more prominent at 9.64 GeV/n, specially for 12C. Furthermore, the structure of such under-populated regions smoothens out with increasing energies until disappearing, a consequence of diffusion being the dominant transport process. In contrast, over-populated regions of particles are present at all energies for secondary species. At 1 TeV the source distribution of the CRs leaves a clear imprint for all species, where secondary CRs can be easily identified by their source distribution being related to the gas distribution.
As a consequence of spallation and energy losses, we see that heavy species such as carbon and boron display more gas-induced structures at the GeV level in contrast to protons and electrons. For primary protons, losses due to interactions with gas are only significant at the densest regions, with inelastic losses and diffusion dominating at the chosen energies and at most locations on the plane (see e.g., Tan and Ng, 1983, for the modeling of the relevant proton cross-section). For electrons, one important process is bremsstrahlung, which dominates electron losses in dense gas regions between 100 MeV and a few GeV, before becoming subdominant to IC and synchroton losses due to interactions with the interstellar radiation field (ISRF) (Evoli et al., 2020). Furthermore, due to the high energy losses at TeV energies, very high energy electrons do not travel far from their production sites – being the ISM gas for secondary electrons – (see e.g., Mertsch, 2020; Evoli et al., 2021; Thaler et al., 2023), leading to a higher contrast and sharper structures in the electron distribution as energy increases, as seen at 1.04 TeV in comparison to 9.64 GeV for both primaries and secondaries.
We now investigate the impact of the gas distribution on CR spectra and the different impact in 2D and 3D models focusing on the B/C ratio. For comparison, we compute all the spectra from the 2D and 3D models shown herein using the same set of transport parameters given in Table 1 for the 3D model. We calculate these spectra at the Galactic center (GC) and at Earth’s location, where the latter is fixed by the underlying velocity field model used to reconstruct the 3D gas density. For the BEG03 model, Earth is assumed to be located at 8.0 kpc from the GC. We chose the GC region to study the impact of the distribution of the gas structures, since the average gas density within any distance from the GC is the same between the 2D and 3D models. For spectra at Earth we compare with CR data, where we use the force-field approximation to model heliospheric modulation. We take the modulation potential (see e.g., Ghelfi et al., 2017, for details on its estimation) from the AMS-02 measurements, as reported in the cosmic ray database (CRDB) (Maurin et al., 2014, 2020, 2023). To illustrate the interplay between primaries and secondaries in each model, in Fig 7 we show the B/C ratio at Earth and at the GC for both the 2D and 3D models. At both locations, as can be seen from the residual plot, we observe a higher B/C ratio in the 2D model than in the 3D model, up to 1 TeV at the GC and extending beyond 1 TeV at Earth. Flux differences at Earth are on average at all energies. At the GC, such differences can be as large as at 300 MeV and decrease with increasing energy. At Earth, the average gas density differs between the 2D and 3D models – with as shown before – leading to differences in the corresponding carbon and boron spectra. For both models, at the GC the local gas density is identical; thus, in this location, spallation and energy losses contribute in the same way to CR fluxes, and the resulting differences can only come as a direct consequence of the spatial distribution of gas structures itself, instead of the local average density. For individual CR species, we also note flux differences. At Earth, we find that CR fluxes below 10 GeV can deviate up to 15% for Carbon, 10% for Boron, 5% for protons, and 2.5% for electrons. At the GC, these differences can double below 10 GeV for all species. We present the spectra of different individual species, including antiprotons, in A.
@@ 3.3 Correlation Structure
Here, we display the correlation structure observed thus far in detail, by focusing on CR protons first to discuss how gas structures modulate their energy-dependent morphology. For this purpose, we study the proton interaction timescales at both the GC and Earth, the low-energy mean distribution of CR protons, and the correlation between gas and CR protons in the 3D model. As previously stated, and shown in A, energy-dependent spatial structures follow from the spatial distribution of gas via the corresponding interactions of the different particle species. When such interactions contribute significantly to CR losses in comparison to the escape rate of CRs from the Galaxy, the resulting CR distribution exhibits local structures that follow the gas maps.
In Fig 8, we show a comparison of different loss processes for protons, i.e. their inelastic, ionization, and escape loss rates averaged over a sphere with radius 4 kpc around Earth. Here, inelastic losses are computed via their inelastic collisions and fragmentation interactions with ISM H and He, following the parametrization in Tan and Ng (1983) for protons. Ionization losses, i.e. the energy loss rate due to the interaction of CRs that traverse neutral matter (i.e. ISM H and He), are adapted from Mannheim and Schlickeiser (1994). Finally, the escape rate is calculated as with kpc the halo height used in our model and the rigidity-dependent diffusion coefficient, using the values for the 3D model in Table 1. Inelastic and ionization losses are directly proportional to the local gas density, while escape rates indirectly depend on the gas distribution by the change of . Given a sufficiently high local gas density, these loss processes drive the presence of under-dense structures in CR primaries, compared to the escape rate of the corresponding CR.
To illustrate the impact of the gas distribution on the distribution of CR fluxes at lower energies, we show in Fig 9 the Galactic plane distribution of primary and secondary protons at 108 MeV/n. Although ionization rates dominate the low energy regime, primary protons are still subject to inelastic losses, which in turn drive the production of secondaries. We note that the ionization losses dominate for energies lower than MeV/n, resulting in decreased primary proton fluxes in most dense gas regions within 10 kpc of the GC.
The under-dense features disappear at locations with sufficiently less dense gas as diffusion becomes dominant, leading to the smoothing of the distribution of primary protons visible in Fig 5 for the distributions 9.64 GeV/n and 1.04 TeV/n. We note that the gas-induced structures that are still visible at 9.64 GeV/n result due to inelastic losses being on a similar magnitude as diffusion. For secondaries, while the loss rates are the same, their injection occurs only at dense gas locations and thus their distribution is primarily determined by the gas maps.
@@
To quantify this energy-dependent relationship between the structure of fluxes for primaries and secondaries, and that of the gas density, we calculate the Pearson correlation coefficient. For this, we use protons and electrons, and compare their distribution with the gas distribution inside 2 regions of radius kpc and kpc around the GC in the Galactic plane. The kpc region considers the small over-dense patch of gas at the very center of the Galaxy, while the kpc region takes into account local over- and under-dense gas clouds. The resulting energy-dependent correlation coefficients, along with the corresponding regions are shown in Fig 10. We note the anti-correlation (correlation) between primaries (secondaries) and gas density in both regions at all energies. For very-high-energy secondary electrons (i.e. TeV/n) the correlation is very close to 1, reflecting an almost linear dependence between gas and secondary electrons as a consequence of their increasingly strong confinement to their source regions (i.e. the high density regions) due to their high energy losses. We observe that with increasing energies, the absolute value of the correlation coefficient between secondaries and gas is bigger than the one between primaries and gas. This comes as a consequence of the anti-correlation with primaries being driven by gas-dependent losses. On the kpc region, the correlation coefficient for secondary protons becomes considerably smaller, since they can travel larger distances than electrons due to lower energy losses.